x/4 - y/11 = -9/44
x/6 - y/13 = -1/26

Answer by stanbon(75887)

x/4 - y/11 = -9/44
x/6 - y/13 = -1/26
-----------------------
Simpify left side of each:
(11x - 4y)/44 = -9/44
(13x - 6y)/78 = -3/78
--------------------------
Equate the numerators:
11x - 4y = -9
13x - 6y = -3
-----------------
Multiply 1st by 3
Multiply 2nd by 2
-------------------------
33x - 12y = -27
26x - 12y = -6
------------------------
Subtract 2nd from 1st to get:
7x = -21
x = -3
-----------------
Substitute into 11x -4y = -9 to solve for "y":
11*-3 -12y = -27
-33 -12y = -27
-12y = 6
y = -1/2
===============
Cheers,
Stan H.